# Two-Level Distributed Interference Management for Large-Scale HAPS-Empowered vHetNets

### v1

@@ Abstract
Next-generation wireless networks (xG) must provide ubiquitous connectivity while enhancing user experience in both densely populated urban areas and rural regions. To achieve this, a disruptive network architecture is essential, and high altitude platform stations (HAPS) offer a promising solution. By integrating HAPS with terrestrial networks, we can create HAPS-empowered vertical heterogeneous networks (vHetNets), which significantly improve coverage and capacity, as well as support emerging use cases. In HAPS-empowered vHetNets, different tiers can share the same spectrum, forming harmonized spectrum vHetNets that enhance spectral efficiency (SE). However, we face two major challenges: i) co-channel interference in harmonized spectrum vHetNets, and ii) the large-scale nature of the network. To address the first challenge, we adopt a cell-free approach as the underlying network architecture for the HAPS-empowered vHetNet. In this approach, base stations use beamforming to direct high-gain, narrow beams toward users, which helps mitigate interference. However, this creates a nonconvex and high-dimensional optimization problem, which highlights the second challenge of dealing with a large-scale network. Consequently, centralized solutions become impractical due to the computational and communication overhead involved. The standard two-block alternating direction method of multipliers (ADMM) is one option, but nonconvex constraints can hinder its convergence. As an alternative, we have developed a two-level distributed proportional fairness beamforming weight design (PFBWD) algorithm. This algorithm uses a combination of the augmented Lagrangian method (ALM) and a three-block ADMM framework. The proposed method effectively tackles nonconvexity, reduces complexity, and enables scalable, distributed optimization with guaranteed convergence.
Index Terms:
HAPS, vHetNet, interference management, cell-free, beamforming, proportional fairness, distributed algorithm
@@ I Introduction
The next-generation (xG) wireless networks are evolving toward AI-driven, sensing-based systems that must deliver immersive throughput, hyper reliability, and low latency, while ensuring connectivity for everyone worldwide [1]. A wide range of innovative technologies is under consideration to meet these goals, spanning multiple layers of the telecommunication network, including terahertz frequency bands, movable antenna architectures, holographic beamforming, semantic communication, and advanced AI-driven resource management strategies. Meeting the ambitious performance and coverage demands requires more than just novel technologies. This also demands innovative network architectures, coupled with efficient performance optimization algorithms that can fully unlock their potential [2]. Among these architectural innovations, non-terrestrial networks (NTN), including space satellites, high altitude platform stations (HAPS), and uncrewed aerial vehicles, represent a particularly promising direction [3]. Unlike earlier deployments that primarily targeted rural and remote areas, future NTN will be tightly integrated with urban terrestrial networks to provide global coverage with seamless quality of experience.
In future xG wireless networks, HAPS is expected to play a critical role, particularly in urban environments with high user density and traffic demand. Positioned quasi-stationary in the stratosphere layer of the Earth at an altitude of approximately 20 km above the ground, HAPS offers several advantages over both space-based satellites and terrestrial networks. For instance, HAPS provides lower latency compared to space satellites, greater flexibility to support advanced technologies (e.g., ultra massive multiple-input multiple-output (umMIMO), reconfigurable intelligent surfaces), and a wider coverage area than terrestrial macro base stations (MBSs). As such, HAPS can serve as a complementary platform with the potential to help meet the demanding requirements of future networks, in conjunction with legacy terrestrial and satellite systems [4]. Moreover, the large surface area of HAPS enables the integration of ambient energy harvesting resources, contributing to the development of a more sustainable network architecture [5, 6]. Referred to by the International Telecommunication Union (ITU) as HAPS as International Mobile Telecommunications (IMT) base stations, HAPS can be integrated with existing terrestrial networks in urban environments to enhance coverage and capacity, thereby supporting the fulfillment of future network KPIs. Such integrated HAPS-terrestrial networks are referred to as HAPS-empowered vertical heterogeneous networks (vHetNets) [7].
@@ I-A Related works
Since the wireless channel is interference-limited, designing efficient interference management strategies has a long research history [10]. Recent studies have explored spectrum sharing and interference suppression techniques in integrated HAPS-terrestrial networks. To this end, the authors in [11] proposed an interference canceler and coordination mechanism to facilitate spectrum sharing between HAPS and terrestrial systems. In [12], the authors introduced a cell design method for HAPS aimed at extending coverage while maintaining coexistence with terrestrial mobile networks. In addition, [13] presented a codebook-based interference suppression approach for space-air-ground integrated networks, demonstrating improved performance in managing inter-tier interference. In our previous work [4], we developed an interference management algorithm for an integrated HAPS-terrestrial network, leveraging subcarrier allocation and power control. Furthermore, in [14] and [15], we formulated three interference management problems aimed at optimizing user association and beamforming weights under different objective functions: weighted sum rate, max-min fairness, and proportional fairness (PF). The performance of the proposed algorithms was subsequently evaluated in both vHetNets and standalone terrestrial network scenarios. More recently, the authors in [16] proposed an interference mitigation framework to enable NTN and terrestrial network coexistence.
@@
Among the existing approaches for managing interference, beamforming weight design has emerged as an efficient strategy [17]. To this end, optimal beamforming weights design algorithms for MIMO systems are developed in a centralized manner [18, 19]. However, due to the high computational and signaling overheads associated with centralized algorithms, significant efforts have also been devoted toward developing distributed algorithms. In  [20], the authors introduced an alternating direction method of multipliers (ADMM)-based beamforming approach for multi-group multicast systems, and in [21], the authors developed a stochastic ADMM framework for coordinated multi-cell beamforming. In addition, the authors in [22] proposed an ADMM-based distributed beamforming algorithm to minimize total transmit power under signal-to-interference-and-noise ratio (SINR) constraints in multi-cell networks. Moreover, [23] applied ADMM to power-efficient downlink beamforming in cell-free massive MIMO, also targeting SINR-constrained optimization across distributed access points.
@@
Although the two-block ADMM approach has been widely adopted in distributed algorithm design [24], its limitations hinder its applicability to general nonconvex problems. To address this, more sophisticated methods have been explored. In [25], the authors formulated a joint user association and beamforming weight design problem aiming at maximizing the ergodic sum rate. Accordingly, the SINR term has been approximated with the average signal-to-leakage-and-noise ratio, and a three-step Gaussian belief propagation-based distributed solver is proposed to solve the problem. In [26], the authors formulated a sum rate maximization problem and introduced a belief propagation algorithm for cooperative MIMO systems that enables efficient downlink beamforming through decentralized message passing. In recent years, deep learning-based approaches have also gained traction. The authors in [27] employ deep reinforcement learning for distributed uplink beamforming in cell-free networks, considering a sum rate objective function. Similarly, [28] uses distributed graph-based learning to jointly optimize user association and beamforming in multi-cell networks to maximize the sum rate. Collectively, these contributions underscore the growing emphasis on distributed beamforming frameworks for xG wireless networks.
@@ I-B Motivations and Contributions
It is evident that centralized algorithms favor small-scale systems; hence, they are practically infeasible in HAPS-empowered vHetNets, which are large-scale systems for several reasons. First, the large coverage area of HAPS encompasses a large number of MBSs, making it a large-scale system. Second, the use of ultra mMIMO involves a large number of antenna elements, resulting in a high-dimensional optimization problem. Therefore, the development of a distributed algorithm for beamforming design becomes necessary for handling large-scale networks such as HAPS-empowered vHetNets. Further, the formulated problems in vHetNets are inherently nonconvex, especially in interference-dominant networks [7]. To address this, various reformulation linearization techniques (RLT) are often employed to approximate the original problem and solve it iteratively through successive convex approximation (SCA) [29]. However, such iterative methods require solving multiple optimization problems iteratively, making them challenging for real-time or low-latency applications in dynamic networks. This reinforces the necessity of developing distributed algorithms tailored for nonconvex interference management problems in vHetNets. That said, directly applying conventional distributed approaches such as two-block ADMM to general nonconvex constrained problems does not guarantee convergence unless relaxation schemes are used [30]. To this end, multiple variants of ADMM approaches have been developed to address the nonconvexity of the optimization problems [31]. In this regard, the authors in [32] proposed a two-level distributed algorithm that can solve the nonconvex constrained problem in a distributed manner, with guaranteed convergence. Specifically, they introduce a key reformulation that guarantees the conditions for convergence of multi-block ADMM algorithms. The resulting two-level algorithm embeds a specially structured three-block ADMM in the inner loop within an augmented Lagrangian method (ALM) framework [32]. This algorithm is selected for its demonstrated advantages in accelerating convergence and improving feasibility in challenging nonconvex settings. Its ability to handle multi-block constraints robustly makes it well-suited for the problem structure.
In this work, we consider a HAPS-empowered vHetNet in which the base stations (BSs) [^1] The term BS in this work collectively refers to HAPS and MBSs. are equipped with MIMO antenna arrays. In [15], we demonstrated the performance gains achieved by associating UEs with multiple BSs. Moreover, the results in [15] showed that the PF objective function offers a fair trade-off between maximizing the sum rate and improving the worst-case UE SE. Therefore, we formulate a cell-free proportional fairness beamforming weight design (PFBWD) problem. The beamforming weights are designed to determine the strength of the signal received by each UE from each BS. Due to the problem’s nonconvex nature, directly applying the ADMM-based optimization approaches doesn’t guarantee convergence. To address this, we reformulate the original problem into an equivalent distributed form and then apply the two-level distributed algorithm. Accordingly, the main contributions of this paper are summarized as follows:
  • We adopt a cell-free approach as the underlying network architecture for HAPS-empowered vHetNet in an urban setting, where the HAPS coverage overlaps with the terrestrial network. A PFBWD problem is formulated to optimize network-wide fairness.
  • We apply the RLT to address the nonconvex nature of the PFBWD problem, and restructure the original problem into an equivalent distributed form.
  • We develop a novel two-level distributed algorithm to solve the large-scale PFBWD problem. The proposed algorithm incorporates a structured three-block ADMM at the inner-level within an outer-level ALM framework.
  • We theoretically establish the conditions required for the convergence of the proposed algorithm.
  • We validate and demonstrate the effectiveness of the proposed two-level distributed algorithm through numerical simulations, comparing it to its centralized counterpart in terms of SE and computational complexity.
[…]
@@ II System Model and Problem Formulation
We consider a cell-free HAPS-empowered vHetNet consisting of one HAPS and $B$ MBSs, simultaneously serving $U$ single antenna UEs as illustrated in Fig. 1. The vHetNet operates in the downlink channel within the sub-6 GHz frequency band. UEs and BSs are indexed by $u\in\mathcal{U}\triangleq\{1,\ldots,U\}$, and $b\in\mathcal{B}\triangleq\{1,\ldots,B+1\}$, respectively, where index $b=B+1$ is reserved for the HAPS. The MBSs and HAPS are equipped with MIMO antennas with a total number of $N_{b}=N^{H}_{b}\times N^{V}_{b}$ antenna elements in a uniform planar array configuration, where $N^{H}_{b}$ and $N^{V}_{b}$ are the number of antenna elements in the horizontal and vertical axes, respectively. We define $\mathbf{H}^{b}=[\mathbf{h}^{b}_{1},\dots,\mathbf{h}^{b}_{U}]\in\mathbb{C}^{N_{b}\times U}$, as the channel coefficients matrix at BS b, where $\mathbf{h}^{b}_{u}=[h^{b}_{1,u},\dots,h^{b}_{N_{b},u}]^{\top}\in\mathbb{C}^{N_{b}}$ denotes the channel vector between BS $b$ and UE $u$. We assume that all UEs share the same time-frequency resources with equal bandwidth allocation. Consequently, interference is propagated from all serving BSs to all UEs. To efficiently manage interference, BSs communicate with UEs using narrow and high-gain beams. Therefore, designing optimal beams is essential and the primary focus of this work.
Due to the strategic positioning of HAPS at an altitude of $20$ km above the ground, UEs experience a different channel compared to that of MBSs. Specifically, the propagation characteristics differ between the HAPS and terrestrial tiers, primarily because of the strong line of sight (LoS) link between HAPS and UEs. As a result, we will consider two distinct channel models, which are explained in detail below.
Figure 1: HAPS-empowered vHetNet system model.

### v2

@@ Abstract
High altitude platform stations (HAPS) offer a promising solution for achieving ubiquitous connectivity in next-generation wireless networks (xG). Integrating HAPS with terrestrial networks, creating HAPS-empowered vertical heterogeneous networks (vHetNets), significantly improves coverage and capacity and supports emerging novel use cases. In HAPS-empowered vHetNets, HAPS and terrestrial network tiers can share the same spectrum, forming harmonized spectrum vHetNets that enhance spectral efficiency (SE). However, harmonized spectrum vHetNets face major challenges, including severe co-channel interference and scalability in large-scale deployments. To address the first challenge, we adopt a cell-free multiple-input multiple-output (MIMO) network architecture in which users are simultaneously served by multiple base stations using beamforming. However, beamforming weight design leads to a nonconvex, high-dimensional optimization problem, highlighting the scalability challenge. To address this second challenge, we develop a two-level distributed proportional fairness beamforming weight design (PFBWD) algorithm. This algorithm combines the augmented Lagrangian method (ALM) with a three-block ADMM framework. Simulation results demonstrate the performance improvements achieved by integrating HAPS with standalone terrestrial networks, as well as the reduced complexity and signaling overhead of the distributed algorithm compared to centralized algorithms.
@@ I Introduction
The next-generation (xG) wireless networks are evolving toward artificial intelligence (AI)-driven, sensing-based systems that must deliver immersive throughput, hyper reliability, and low latency, while ensuring connectivity for everyone worldwide [1]. A wide range of innovative technologies is under consideration to meet these goals, spanning multiple layers of the telecommunication network, including terahertz frequency bands, movable antenna architectures, holographic beamforming, semantic communication, and advanced AI-driven resource management strategies. Meeting the ambitious performance and coverage demands requires more than just novel technologies. This also demands innovative network architectures, coupled with efficient performance optimization algorithms that can fully unlock their potential [2]. Among these architectural innovations, non-terrestrial networks (NTN), including space satellites, high altitude platform stations (HAPS), and uncrewed aerial vehicles, represent a particularly promising direction [3]. Unlike earlier deployments that primarily targeted rural and remote areas, future NTN will be tightly integrated with urban terrestrial networks to provide global coverage with seamless quality of experience.
In future xG wireless networks, HAPS is expected to play a critical role, particularly in urban environments with high user density and traffic demand. Positioned quasi-stationary in the stratosphere layer of the Earth at an altitude of approximately 20 km above the ground, HAPS offers several advantages over both space-based satellites and terrestrial networks. For instance, HAPS provides lower latency compared to space satellites, greater flexibility to support advanced technologies (e.g., ultra massive multiple-input multiple-output (umMIMO), reconfigurable intelligent surfaces), and a wider coverage area than terrestrial macro base stations (MBSs). As such, HAPS can serve as a complementary platform with the potential to help meet the demanding requirements of future networks, in conjunction with legacy terrestrial and satellite systems [4]. Moreover, the large surface area of HAPS enables the integration of ambient energy harvesting resources, contributing to the development of a more sustainable network architecture [5, 6]. Referred to by the International Telecommunication Union (ITU) as HAPS as International Mobile Telecommunications (IMT) base stations (HIBS), HAPS can be integrated with existing terrestrial networks in urban environments to enhance coverage and capacity, thereby supporting the fulfillment of future network KPIs such as ubiquitous connectivity, immersive data rate, and sustainability. Such integrated HAPS-terrestrial networks are referred to as HAPS-empowered vertical heterogeneous networks (vHetNets) [7].
@@ I-A Related works
Since the wireless channel is interference-limited, designing efficient interference management strategies has a long research history [10]. Recent studies have explored spectrum sharing and interference suppression techniques in integrated HAPS-terrestrial networks. To this end, the authors in [11] proposed an interference canceler and coordination mechanism to facilitate spectrum sharing between HAPS and terrestrial systems. In [12], the authors introduced a cell design method for HAPS aimed at extending coverage while maintaining coexistence with terrestrial mobile networks. In addition, [13] presented a codebook-based interference suppression approach for space-air-ground integrated networks, demonstrating improved performance in managing inter-tier interference. In our previous work [4], we developed an interference management algorithm for an integrated HAPS-terrestrial network, leveraging subcarrier allocation and power control. More recently, the authors in [14] proposed an interference mitigation framework to enable coexistence between NTN and terrestrial networks.
Among the existing approaches for managing interference, beamforming weight design has emerged as an efficient strategy [15]. To this end, in [16] and [17], we developed three centralized interference management algorithms aimed at optimizing user association and beamforming weights under different objective functions: weighted sum rate (WSR), max-min fairness (MMF), and proportional fairness (PF). The performance of the proposed algorithms was subsequently evaluated in both vHetNets and standalone terrestrial network scenarios. Furthermore, the authors in [18] formulated a max-min sum rate beamforming design problem in a HetNet with hardware impairments. In [19], a centralized beamforming weight design algorithm was proposed for cell-free terrestrial networks to optimize the tradeoff between energy efficiency and SE. However, due to the high computational and signaling overheads associated with centralized algorithms, significant efforts have also been devoted toward developing distributed algorithms.
@@
Accordingly, in [20], the authors introduced an alternating direction method of multipliers (ADMM)-based beamforming approach for multi-group multicast systems under two objective functions of minimizing the total transmit power and MMF. In [21], the authors developed a stochastic ADMM framework for coordinated multi-cell beamforming in a smartgrid powered coordinated multicell downlink system. In addition, the authors in [22] proposed an ADMM-based distributed beamforming algorithm to minimize total transmit power under signal-to-interference-and-noise ratio (SINR) constraints in multi-cell networks. Moreover, [23] applied ADMM to power-efficient downlink beamforming in cell-free mMIMO, also targeting SINR-constrained optimization across distributed access points. In [24], the authors proposed a deep reinforcement learning (DRL)-based beamforming approach for mMIMO aerial communications, aiming at maximizing the sum rate.
@@
Although the two-block ADMM approach has been widely adopted in distributed algorithm design [25], its limitations hinder its applicability to general nonconvex problems. To address this, more sophisticated methods have been explored. In [26], the authors formulated a joint user association and beamforming weight design problem aiming at maximizing the ergodic sum rate. Accordingly, the SINR term has been approximated with the average signal-to-leakage-and-noise ratio, and a three-step Gaussian belief propagation-based distributed solver is proposed to solve the problem. In [27], the authors formulated a sum rate maximization problem and introduced a belief propagation algorithm for cooperative MIMO systems that enables efficient downlink beamforming through decentralized message passing. In recent years, deep learning-based approaches have also gained traction. The authors in [28] employ DRL for distributed uplink beamforming in cell-free networks, considering a sum rate objective function. Similarly, [29] uses distributed graph-based learning to jointly optimize user association and beamforming in multi-cell networks to maximize the sum rate. Collectively, these contributions underscore the growing emphasis on distributed beamforming frameworks for xG wireless networks. Existing distributed studies are primarily developed for terrestrial networks operating at relatively smaller scales compared to HAPS-empowered vHetNets. These works typically focus on sum rate maximization, which often prioritizes overall throughput at the expense of the minimum achievable SE. Moreover, to address the inherent nonconvexity of the formulated optimization problems, existing approaches rely on simplifying approximations or relaxations, which can limit the achievable performance. Consequently, distributed solutions that jointly address scalability, fairness guarantees, and harmonized-spectrum operation in cell-free HAPS-empowered vHetNets remain largely unexplored.
@@ I-B Motivations and Contributions
It is evident that centralized algorithms favor small-scale systems; hence, they are practically infeasible in harmonized spectrum HAPS-empowered vHetNets, which are large-scale systems for several reasons. First, the large coverage area of HAPS encompasses a large number of MBSs, serving a large number of UEs. Since HAPS and MBSs are sharing the same spectrum, they need to coordinate with each other in order to serve UEs. Considering the large coverage area of HAPS, this creates a large-scale network. Second, the use of ultra mMIMO involves a large number of antenna elements, particularly at HAPS with a large surface area, resulting in a high-dimensional optimization problem. Therefore, the development of a distributed algorithm for beamforming design becomes necessary for handling large-scale networks such as HAPS-empowered vHetNets. Furthermore, centralized algorithms must collect all network information, including channel coefficients, solve the optimization problem, and distribute the results to the base stations (BSs [^1] The term BS in this work collectively refers to HAPS and MBSs.). However, exchanging data between HAPS and MBSs, considering the positioning of HAPS, imposes additional latency to the network. Therefore, distributed algorithms to design the important network parameters are a key solution to overcome these challenges.
@@
In addition to the aforementioned challenges, the formulated problems in vHetNets are inherently nonconvex, especially in interference-dominant networks [7]. To address this, various reformulation linearization techniques (RLT) are often employed to approximate the original problem and solve it iteratively through successive convex approximation (SCA) [30]. However, such iterative methods require solving multiple optimization problems iteratively, making them challenging for real-time or low-latency applications in dynamic networks. This reinforces the necessity of developing distributed algorithms tailored for nonconvex interference management problems in vHetNets. However, directly applying conventional distributed approaches such as two-block ADMM to general nonconvex constrained problems does not guarantee convergence unless relaxation schemes are used [31]. To this end, multiple variants of ADMM approaches have been developed to address the nonconvexity of the optimization problems [32]. Accordingly, the authors in [33] proposed a two-level distributed algorithm that can solve the nonconvex constrained problem in a distributed manner, with guaranteed convergence. Specifically, they introduce a key reformulation that guarantees the conditions for convergence of multi-block ADMM algorithms. The resulting two-level algorithm embeds a specially structured three-block ADMM in the inner loop within an augmented Lagrangian method (ALM) framework [33]. This algorithm is selected for its demonstrated advantages in accelerating convergence and improving feasibility in challenging nonconvex settings. Its ability to handle multi-block constraints robustly makes it well-suited for the problem structure.
In this work, we consider a harmonized spectrum HAPS-empowered vHetNet, where HAPS and MBSs are serving UEs in an overlapped coverage area. This novel network architecture is a potential solution to improve coverage and capacity in urban areas. In the system model, the BSs are equipped with MIMO antenna arrays, empowering the BSs to serve UEs using beamforming. Our earlier works investigated related but fundamentally different system models. In [16], a centralized algorithm was developed for joint user association and beamforming, where each UE was served by only one BS (either the HAPS or an MBS). The follow-up study in [17] extended the analysis to multiple objective functions (MMF, PF, and WSR) but still relied on a centralized optimization framework and a single-BS association model. These studies revealed two important insights: (i) allowing UEs to be served by multiple BSs can significantly improve spectral efficiency [16], and (ii) PF provides a desirable balance between maximizing the sum data rate and the minimum SE of the network. Motivated by these observations, in this paper, we adopt a cell-free architecture in which UEs are jointly served by MBSs and HAPS, simultaneously. Despite recent advances in cell-free architectures and distributed optimization, how to enforce proportional fairness in harmonized-spectrum vHetNets under severe inter-tier interference, without centralized CSI aggregation, remains an open problem.
We formulate a PF beamforming weight design (PFBWD) problem tailored to this architecture. Unlike [16, 17], which rely on centralized optimization and single-BS association, this work addresses a cell-free HAPS-empowered vHetNet architecture under a harmonized spectrum, where centralized CSI collection is infeasible. Moreover, while [33] proposes a generic two-level distributed framework, this paper makes a nontrivial contribution by reformulating a highly coupled PFBWD problem in HAPS-empowered vHetNets into a structure that satisfies the convergence conditions of [33], which is not straightforward due to the interference coupling and network-side fairness objective function. To this end, we reformulate the original problem into an equivalent distributed structure and develop a two-level distributed algorithm in which each BS optimizes its local variables using only local channel coefficients, without exchanging channel coefficients.
Accordingly, the main contributions of this paper are summarized as follows:
  • We study a harmonized spectrum HAPS-empowered vHetNet with overlapping HAPS–terrestrial coverage under a cell-free architecture, where UEs are jointly served by multiple MBSs and the HAPS. In this setting, we formulate a PFBWD optimization problem that explicitly captures the strong inter-tier interference coupling inherent to such networks.
  • We show that the resulting PFBWD problem cannot be efficiently solved using centralized optimization due to its large scale, dense interference structure, and the impracticality of global channel coefficient collection in HAPS-empowered vHetNets. To address this, we reformulate the problem into an equivalent distributed structure that enables BS-level optimization using only local channel state information, without inter-BS channel exchange.
  • Based on the proposed reformulation, we develop a two-level distributed optimization framework in which a structured three-block ADMM is embedded within an outer ALM. The proposed framework is tailored to the considered cell-free HAPS-empowered vHetNet and is capable of handling the nonconvex formulated PFBWD problem.
  • We establish the convergence of the proposed two-level distributed algorithm and validate its performance for large-scale nonconvex interference management through simulations.
Nomenclature: We use italic lowercase letter $x$ for scalars, bold lowercase letter $\mathbf{x}$ for vectors, bold uppercase letter $\mathbf{X}$ for matrices, and calligraphic uppercase letter $\mathcal{X}$ for sets. The $i$th entry of vector $\mathbf{x}$ is denoted by $\mathbf{x}_{i}$, and the $(i,j)$ entry of matrix $\mathbf{X}$ by $\mathbf{X}_{i,j}$. The notation $\|\cdot\|$ represents the Euclidean norm, while $\|\cdot\|_{F}$ denotes the Frobenius norm. The Kronecker product is represented by $\otimes$. The superscript $(\cdot)^{\top}$ and $(\cdot)^{H}$ denote transpose and Hermitian transpose, respectively.
@@
Figure 1: HAPS-empowered vHetNet system model.
@@ II System Model and Problem Formulation
We consider a cell-free HAPS-empowered vHetNet consisting of one HAPS and $B$ MBSs, simultaneously serving $U$ single antenna UEs as illustrated in Fig. 1. The vHetNet operates in the downlink channel within the sub-6 GHz frequency band. UEs and BSs are indexed by $u\in\mathcal{U}\triangleq\{1,\ldots,U\}$, and $b\in\mathcal{B}\triangleq\{1,\ldots,B+1\}$, respectively, where index $b=B+1$ is reserved for the HAPS. The MBSs and HAPS are equipped with MIMO antennas with a total number of $N_{b}=N^{H}_{b}\times N^{V}_{b}$ antenna elements in a uniform planar array configuration, where $N^{H}_{b}$ and $N^{V}_{b}$ are the number of antenna elements in the horizontal and vertical axes, respectively. To efficiently manage interference, BSs employ beamforming to communicate with UEs using user-specific beams. Therefore, the proper design of beamforming weights becomes essential and is the primary focus of this work. We denote the beamforming weight matrix at BS $b$ by $\mathbf{W}^{b}=[\mathbf{w}^{b}_{1},\dots,\mathbf{w}^{b}_{U}]\in\mathbb{C}^{N_{b}\times U}$, where $\mathbf{w}^{b}_{u}\in\mathbb{C}^{N_{b}}$ refers to the beamforming vector for UE $u$ at BS $b$, and each element $w^{b}_{r,u}$ represents the complex weight applied to antenna element $r\in\{1,\dots,N_{b}\}$ when transmitting to UE $u$. Accordingly, we define $\mathbf{H}^{b}=[\mathbf{h}^{b}_{1},\dots,\mathbf{h}^{b}_{U}]\in\mathbb{C}^{N_{b}\times U}$, as the channel coefficients matrix at BS b, where $\mathbf{h}^{b}_{u}=[h^{b}_{1,u},\dots,h^{b}_{N_{b},u}]^{\top}\in\mathbb{C}^{N_{b}}$ denotes the channel vector between BS $b$ and UE $u$. We assume that all UEs share the same time-frequency resources with equal bandwidth allocation. Consequently, interference is propagated from all serving BSs to all UEs.
Due to the strategic positioning of the HAPS at an altitude of $20$ km above the ground, its role in the network and its interaction with MBSs and UEs are fundamentally different from those of terrestrial BSs, including both MBSs and small-cell BSs (SBSs). In particular, the large elevation angle and strong line-of-sight (LoS) connection between the HAPS and a large number of serving UEs result in channel characteristics that differ significantly from terrestrial links, which are typically dominated by non-LoS (NLoS) propagation and localized interference. As a consequence, HAPS transmissions generate a broad interference footprint that simultaneously impacts a large number of UEs and MBSs over an extended area, necessitating large-scale coordination under harmonized spectrum operation. This behavior is fundamentally different from conventional MBS–SBS deployments, where interference is spatially confined and coordination is typically limited to a small cluster of neighboring cells. Accordingly, we adopt two distinct channel models for the HAPS and terrestrial tiers, which are described in detail in the following subsection.